# Exact electromagnetic multipole expansion using elementary current multipoles

### v1

@@ Conclusions
where $r$, $\theta$, and $\phi$ are the spherical coordinates and $\tilde{M}^{(l)}(\hat{\textbf{v}},m)$ is defined as
$\tilde{M}^{(l)}(\hat{\textbf{v}},m)=\int J_{v}(\textbf{r})j_{l-1}(kr)Y_{l-1,m}^{*}(\theta,\phi)d^{3}\textbf{r}.$ (19)
Let us find several lowest-order coefficients $\tilde{M}^{(l)}(\hat{\textbf{v}},m)$ with the spherical harmonics written in Cartesian coordinates:
$\tilde{M}^{(1)}(\hat{\textbf{v}},0)=\frac{1}{2}\sqrt{\frac{1}{\pi}}\int J_{v}(\textbf{r})j_{0}(kr)d^{3}\textbf{r},$ (20)
$\tilde{M}^{(2)}(\hat{\textbf{v}},0)=\frac{1}{2}\sqrt{\frac{3}{\pi}}\int J_{v}(\textbf{r})z\frac{j_{1}(kr)}{r}d^{3}\textbf{r},$
$\tilde{M}^{(2)}(\hat{\textbf{v}},\pm 1)\!=\!\mp\frac{1}{2}\sqrt{\frac{3}{2\pi}}\int J_{v}(\textbf{r})(x\pm iy)\frac{j_{1}(kr)}{r}d^{3}\textbf{r},$
$\tilde{M}^{(3)}(\hat{\textbf{v}},0)=\frac{1}{4}\sqrt{\frac{5}{\pi}}\int J_{v}(\textbf{r})(3z^{2}-r^{2})\frac{j_{2}(kr)}{r^{2}}d^{3}\textbf{r},$
$\tilde{M}^{(3)}(\hat{\textbf{v}},\pm 1)=\mp\frac{1}{2}\sqrt{\frac{15}{2\pi}}\int J_{v}(\textbf{r})(xz\pm iyz)\frac{j_{2}(kr)}{r^{2}}d^{3}\textbf{r}.$
Each $\tilde{M}^{(l)}(\hat{\textbf{v}},m)$ can be written as a linear superposition of the following normalized multipole moments:
$M^{(l)}(\hat{\textbf{v}},a,b)=$ $\,\frac{i}{\omega}\frac{(2l-1)!!}{(l-1)!}\int J_{v}(\textbf{r})$ (21)
@@
These moments form the general current multipole tensor defined in Eq. (3). In the limit of $kr\ll 1$, $j_{l-1}(kr)/(kr)^{l-1}$ approaches 1/(2l1)!!1/(2l-1)!! [33], and Eq. (21) converges to Eq. (14). Correspondingly, Eq. (3) converges to Eq. (2).
The expression for the scattering current density in terms of the exact current multipole moments can be obtained from Eq. (12). Expanding the Dirac delta function in this equation in terms of spherical harmonics, we find:
$\textbf{J}(\textbf{r})=$ $\,\frac{k}{\pi}\sum_{l=1}^{\infty}\,\sum_{\hat{\textbf{v}}=\hat{\textbf{x}},\hat{\textbf{y}},\hat{\textbf{z}}}\,\sum_{m=-l+1}^{l-1}\frac{2l-1}{r^{2}}j_{l-1}(kr)$ (22)
where each coefficient $\tilde{M}^{(l)}(\hat{\textbf{v}},m)$ is written in terms of multipole moments $M^{(l)}(\hat{\textbf{v}},a,b)$. The contribution of each current multipole to the near field of a scatterer can be assessed either from Eq. (18) or from Eq. (22), since the electric field is proportional to both A(r) and J(r). For example, if the characteristic size of the scatterer $L$ satisfies the condition $kL\ll 1$, the contributions of the current dipole, quadrupole, and octupole scale proportionally to $\textbf{{D}}/L$, $\textbf{{Q}}/L^{2}$, and $\textbf{{O}}/L^{3}$, respectively, since
$h_{0}^{(1)}(kL)=-ie^{ikL}\frac{1}{kL},$
$h_{1}^{(1)}(kL)=-e^{ikL}\frac{kL+i}{(kL)^{2}},$
$h_{2}^{(1)}(kL)=ie^{ikL}\frac{(kL)^{2}+3ikL-3}{(kL)^{3}}.$ (23)
Functions $j_{l-1}(kL)$ scale similarly. For larger scatterers, all the terms in the expressions for $h_{l-1}^{(1)}(kL)$ or $j_{l-1}(kL)$ must be included.
Appendix B Classical multipole moments as linear superpositions of current multipole moments.

### v2

@@ Conclusions
where $r$, $\theta$, and $\phi$ are the spherical coordinates, and the coefficients $\tilde{M}^{(l)}(\hat{\textbf{v}},m)$ are given by
$\tilde{M}^{(l)}(\hat{\textbf{v}},m)=\int\limits_{-\infty}^{\infty}J_{v}(\textbf{r})\frac{j_{l-1}(kr)}{k^{l-1}}Y_{l-1,m}^{*}(\theta,\phi)d^{3}\textbf{r}.$ (40)
These coefficients are the current multipole moments in the expansion of $\textbf{A}(\textbf{r})$. They possess mutually orthogonal current distributions and can be used as a basis for expansion of arbitrary localized current densities. Let us find several lowest-order moments $\tilde{M}^{(l)}(\hat{\textbf{v}},m)$ with the spherical harmonics in their expressions written in Cartesian coordinates:
$\tilde{M}^{(1)}(\hat{\textbf{v}},0)=\frac{1}{2}\sqrt{\frac{1}{\pi}}\int\limits_{-\infty}^{\infty}J_{v}(\textbf{r})j_{0}(kr)d^{3}\textbf{r},$ (41)
$\tilde{M}^{(2)}(\hat{\textbf{v}},0)=\frac{1}{2}\sqrt{\frac{3}{\pi}}\int\limits_{-\infty}^{\infty}J_{v}(\textbf{r})z\frac{j_{1}(kr)}{kr}d^{3}\textbf{r},$
$\tilde{M}^{(2)}(\hat{\textbf{v}},\pm 1)\!=\!\mp\frac{1}{2}\sqrt{\frac{3}{2\pi}}\int\limits_{-\infty}^{\infty}J_{v}(\textbf{r})(x\pm iy)\frac{j_{1}(kr)}{kr}d^{3}\textbf{r},$
$\tilde{M}^{(3)}(\hat{\textbf{v}},0)=\frac{1}{4}\sqrt{\frac{5}{\pi}}\int\limits_{-\infty}^{\infty}J_{v}(\textbf{r})(3z^{2}-r^{2})\frac{j_{2}(kr)}{(kr)^{2}}d^{3}\textbf{r},$
$\tilde{M}^{(3)}(\hat{\textbf{v}},\pm 1)=\mp\frac{1}{2}\sqrt{\frac{15}{2\pi}}\int\limits_{-\infty}^{\infty}J_{v}(\textbf{r})(xz\pm iyz)\frac{j_{2}(kr)}{(kr)^{2}}d^{3}\textbf{r}.$
It can be seen that each moment $\tilde{M}^{(l)}(\hat{\textbf{v}},m)$ is a simple linear combination of the following normalized moments:
$M^{(l)}(\hat{\textbf{v}},a,b)=$ $\,\frac{i}{\omega}\frac{(2l-1)!!}{(l-1)!}\int\limits_{-\infty}^{\infty}J_{v}(\textbf{r})$ (42)
@@
These moments form the general scattering-current multipole tensor defined in Eq. (3). In the limit of $kr\ll 1$, $j_{l-1}(kr)/(kr)^{l-1}$ approaches 1/(2l1)!!1/(2l-1)!! [28], and Eq. (42) converges to Eq. (32). Correspondingly, Eq. (3) converges to Eq. (2).
Appendix B Current density associated with current multipoles.
For a point-like scatterer, Eq. (35) connects the current multipole moments with the associated scattering current densities. For a realistic scatterer that is much smaller than the wavelength, the current density in a dipole excitation is nearly uniform. Hence, the volume-averaged current density can be equated to the current density itself, and we have
$J^{(1)}_{\hat{\textbf{v}},0,0}(\textbf{r})=-i\omega p_{v}\delta(\textbf{r})=-i\omega p_{v}\frac{1}{V},$ (43)
where $V$ is the volume of the scatterer. Then, to keep the relative strengths of the multipolar current densities unchanged, we replace $\delta(\textbf{r})$ with $1/V$ also for other orders $l$ in Eq. (35). Furthermore, in the considered model, the current density is required to produce the actual field only outside the scatterer, and the field and current density inside the scatterer can be arbitrary. Therefore, we expect the calculated current density to correspond to the realistic current density only on the surface of the scatterer. If the radial sizes of the scatterer along the $x$, $y$, and $z$ directions are $X,Y$, and $Z$, respectively (such that the volume of the scatterer is $V\approx 8XYZ$), the approximate expression for the multipole current density on the surface can be written as
$J^{(l)}(\hat{\textbf{v}},a,b)\approx-\frac{i}{8}\frac{\omega(l-1)!}{X^{a+1}Y^{b+1}Z^{l-a-b}}M^{(l)}(\hat{\textbf{v}},a,b).$ (44)
This equation allows assessing the surface current density and the corresponding near-field enhancement associated with a given multipole moment. The phase of $J^{(l)}(\hat{\textbf{v}},a,b)$ is equal to that of the outermost current element in the given multipole at positive $x$, $y$, and $z$.
The generalized surface current density associated with each multipole moment can be written in the same form as Eq. (44), but with $V=8XYZ$ being replaced with
$I=(2l-1)!!\int\limits_{V}\frac{j_{l-1}(kr)}{(kr)^{l-1}}d^{3}\textbf{r}.$ (45)
Here, the integration is performed over the volume of the scatterer. This gives the following result:
$J^{(l)}(\hat{\textbf{v}},a,b)\approx-\frac{i}{I}\frac{\omega(l-1)!}{X^{a}Y^{b}Z^{l-a-b-1}}M^{(l)}(\hat{\textbf{v}},a,b).$ (46)
If the scatterer is much smaller than the wavelength, $I$ is equal to $V$, which makes Eq. (46) identical to Eq. (44).
Appendix C Classical multipole moments as linear superpositions of current multipole moments.